McCain Foods has a market value equal to its book value. Currently, the firm has excess cash of $9,156 and other assets of $74,200. Equity is worth $45,300. The firm has 1,450 shares of stock outstanding and net income of $5,800. What will the new earnings per share be if the firm uses its excess cash to complete a stock repurchase?

$5.94

$6.29

$5.01

$5.77

$5.42
Answer #1
Equity value $45,300
Shares outstanding pre repurchase 1450
Price per share                                 $45,300/1450 = $31.2414
(Equity value/Shares outstanding)
Excess Cash $9,156
No. of share company can repurchase 9156/31.2414 = 293
(excess cash / price per share)
Shares outstanding post repurchase 1450 - 293 = 1157
(no. shares pre repurchase - no shares bought back )
Net Income $5,800
Earning per share post repurchase $5,800/1157 = $5.01
(Net income / no. of shares outstanding post repurchase)